# AI-enhanced tuning of quantum dot Hamiltonians toward Majorana modes

### v1

@@ VI Methods
with the tunneling matrix defined as $W=\mathrm{diag}(1,0,1)\otimes\mathrm{diag}(\sqrt{\Gamma},\sqrt{\Gamma},-\sqrt{\Gamma},-\sqrt{\Gamma})$ and the dot-lead coupling $\Gamma=0.1$ meV. If we reshape the $S$-matrix accordingly $S=S_{n,p,s,n^{\prime}\!,p^{\prime}\!,s^{\prime}}$ ($n=L,C,R$ indexing dots, $p=1,2$ particle, and $s$ – spin subspaces) then the respective reflection matrices: $s^{ee}_{s,s^{\prime}}\!(i,j)=S_{i,1,s,j,1,s^{\prime}}$ and $s^{he}_{s,s^{\prime}}\!(i,j)=S_{i,2,s,j,1,s^{\prime}}$ give differential conductance as, $ss^{\dagger}\equiv|s|^{2}$,
@@
in unit of $e^{2}\!/h$, $i,j=L,R$ denoting left ($L$) or right ($R$) lead, and $E_{F}$ being the Fermi energy in the leads. The input maps include 4 conductance $G_{ij}$ components: $G_{\mathrm{LL}}$, $G_{\mathrm{LR}}$, $G_{\mathrm{RL}}$, and $G_{\mathrm{RR}}$, where for instance $G_{\mathrm{LL}}=\frac{dI_{\mathrm{L}}}{dV_{\mathrm{L}}}$, with $I_{\mathrm{L}}$ denoting current through the left lead, and $V_{\mathrm{L}}$ is the bias voltage of the left lead. Similarly, other components can be defined by using different combinations of left (L) and right (R) leads. Noteworthy, each conductance map is defined by 2D plot of $G_{ij}$ as a function of some parameter and $E_{F}$. We utilize 4 maps per each component: 3 for $\mu_{i=\mathrm{L},\mathrm{C},\mathrm{R}}$ variation and 1 for $V_{\mathrm{Z}}$ variation: 16 maps in total serving as the input. Fig. 6 shows the conductance maps ($G_{LL}$ component) for the reference parameters $P_{0}$, highlighting the emerged zero-biased peaks.
@@
with $\{\psi_{i},E_{i}\}$ being the collection of $H(P)$ eigenpairs. Both $m_{i}$ and $p_{i}$ are sorted in ascending order by $|E_{i}|$. Eigenstates projection $m_{i}$ on the left (or equivalently right) Majorana mode $M=\gamma_{L}$ is weighted by their distance from zero energy $|E_{i}|$ with some threshold $\epsilon=0.1$ meV. Second factor – $p_{i}$ quantifies electron-hole symmetry: $u_{i},v_{i}$ are the electron-hole components of eigenvectors $\psi_{i}$. By the $H$ symmetry $m_{0}=m_{1}$, and also $p_{0}=p_{1}$. For ideal MZM pair we should have $m_{0}=m_{1}=1$ and $p_{0}=p_{1}=1$, giving $\mathcal{M}=1$. The form of the metric, Eq. 5, is designed to discriminate trivial zero modes described e.g. in the SI of Ref. [Bordin2025], especially $\sum_{i>1}m_{i}$ term penalizes the presence of trivial zero modes, even when the topological ones are there as well.
In our work we adapt typical ViT architecture [dosovitskiy2021] adjusted to process 16-channel $50\times 50$ input tensor (representing all generated conductance map components) and returning vector of the Hamiltonian parameter corrections $\delta P$. The architecture hyperparameters such as number of attention heads (4), number of attention blocks (6), hidden size (256) and patch size (1) are tuned to decrease value of $\mathcal{L}$ after 100 epochs below given threshold $\tau=-0.2$.

### v3

@@ V Methods
with the tunneling matrix defined as $W=\mathrm{diag}(1,0,1)\otimes\mathrm{diag}(\sqrt{\Gamma},\sqrt{\Gamma},-\sqrt{\Gamma},-\sqrt{\Gamma})$ and the dot-lead coupling $\Gamma=0.1$ meV. If we reshape the $S$-matrix accordingly $S=S_{n,p,s,n^{\prime}\!,p^{\prime}\!,s^{\prime}}$ ($n=\mathrm{L,C,R}$ indexing dots, $p=1,2$ particle, and $s$ – spin subspaces) then the respective reflection matrices: $s^{ee}_{s,s^{\prime}}\!(i,j)=S_{i,1,s,j,1,s^{\prime}}$ and $s^{he}_{s,s^{\prime}}\!(i,j)=S_{i,2,s,j,1,s^{\prime}}$ give differential conductance as, $ss^{\dagger}\equiv|s|^{2}$,
@@
in unit of $e^{2}\!/h$, $i,j=\mathrm{L,R}$ denoting left (L) or right (R) lead, and $E_{F}$ being the Fermi energy in the leads. The input maps include 4 conductance $G_{ij}$ components: $G_{\mathrm{LL}}$, $G_{\mathrm{LR}}$, $G_{\mathrm{RL}}$, and $G_{\mathrm{RR}}$, where for instance $G_{\mathrm{LL}}=\frac{dI_{\mathrm{L}}}{dV_{\mathrm{L}}}$, with $I_{\mathrm{L}}$ denoting current through the left lead, and $V_{\mathrm{L}}$ is the bias voltage of the left lead. Similarly, other components can be defined by using different combinations of left (L) and right (R) leads. Noteworthy, each conductance map is defined by 2D plot of $G_{ij}$ as a function of some parameter and $E_{F}$. We utilize 4 maps per each component: 3 for $\mu_{i=\mathrm{L},\mathrm{C},\mathrm{R}}$ variation and 1 for $V_{\mathrm{Z}}$ variation: 16 maps in total serving as the input. Fig. 6 shows the conductance maps ($G_{\mathrm{LL}}$ component) for the reference parameters $P_{0}$, highlighting the emerged zero-biased peaks.
@@
with $\{\psi_{i},E_{i}\}$ being the collection of $H(P)$ eigenpairs. The index $i$ enumerates the subsequent eigenvalues: both $m_{i}$ and $p_{i}$ are sorted in ascending order by $|E_{i}|$. Eigenstates projection $m_{i}$ on the left (or equivalently right) Majorana mode $M=\gamma_{\mathrm{L}}$ is weighted by their distance from zero energy $|E_{i}|$ with some threshold $\epsilon=0.1$ meV. Second factor – $p_{i}$ quantifies electron-hole symmetry: $u_{i},v_{i}$ are the electron-hole components of eigenvectors $\psi_{i}$. By the $H$ symmetry $m_{0}=m_{1}$, and also $p_{0}=p_{1}$. For ideal MZM pair we should have $m_{0}=m_{1}=1$ and $p_{0}=p_{1}=1$, giving $\mathcal{M}=1$. The form of the metric, Eq. 5, is designed to discriminate trivial zero modes such as ABS, which discrimination from MZMs is described e.g. in Ref. [5, 26, 37]. Especially, the $\sum_{i>1}m_{i}$ term penalizes the presence of trivial zero modes, even when the topological ones are there as well.
Although the metric defined in Eq. (5) contains piecewise operations such as $\max(0,\cdot)$ and sorting by $|E_{i}|$, gradients required for training are computed using automatic differentiation provided by the autograd framework. Such piecewise functions are handled using standard subgradient rules (analogous to ReLU activations), so that they propagate whenever the argument of the $\max(0,\cdot)$ operation is positive. The eigenvalues and eigenvectors are obtained from a differentiable eigensolver applied to the Hamiltonian. Degeneracies and level crossings occur only on sets of measure zero in parameter space and did not cause practical issues during training. However, eigenvector-based gradients may become numerically ill-conditioned near degeneracies or for larger matrices, which can introduce noise in the optimization signal. This potential limitation is discussed further in Appendix C.
For the visualization of the spectra in Figs. 5 and 7, two auxiliary quantities are computed from the normalized BdG eigenvectors $\psi_{i}$, utilizing their electron and hole components $u_{i}$, $v_{i}$. The site occupation of dot $n$ for eigenstate $i$ is defined as
$O^{(i)}_{n}=|u_{i,n}|^{2}+|v_{i,n}|^{2}.$ (6)
The quantity referred to as edge occupations is the total weight on the outer dots,
$O_{edge}^{(i)}=O^{(i)}_{\mathrm{L}}+O^{(i)}_{\mathrm{R}}$. (7)
In addition, electron-hole symmetry is defined as the difference between the total electron and hole weights
$S_{\mathrm{eh}}^{(i)}=\sum_{n}\left(|u_{i,n}|^{2}-|v_{i,n}|^{2}\right).$ (8)
In our work we adapt typical ViT architecture [11] adjusted to process 16-channel $50\times 50$ input tensor (representing all generated conductance map components) and returning vector of the Hamiltonian parameter corrections $\delta P$. The architecture hyperparameters such as number of attention heads (4), number of attention blocks (6), hidden size (256) and patch size (1) are tuned to decrease value of $\mathcal{L}$ after 100 epochs below given threshold $\tau=-0.7$.
@@ Acknowledgements
Appendix A Appendix A: Analytical considerations for 3 QDs model
Let’s start with the Rashba-Zeeman-BCS Hamiltonian, Eq. 1, written in so-called Nambu spinor representation $\hat{\Psi}_{n}=(c_{n\uparrow},c_{n\downarrow},c^{\dagger}_{n\uparrow},c^{\dagger}_{n\downarrow})^{T}$: $H=\frac{1}{2}\sum_{nm}\hat{\Psi}^{\dagger}_{n}H^{\mathrm{BdG}}_{nm}\hat{\Psi}_{m}$, where $H^{\mathrm{BdG}}$ is a first-quantized Hamiltonian, also called the Bogoliubov-de Gennes (BdG) Hamiltonian. It’s explicit matrix form reads:
$H^{\mathrm{BdG}}=$
$\sum_{n=1}^{N}\begin{pmatrix}-\mu_{n}+V_{\mathrm{Z}}&0&0&\Delta_{n}e^{i\phi_{n}}\\ 0&-\mu_{n}-V_{\mathrm{Z}}&-\Delta_{n}e^{i\phi_{n}}&0\\ 0&-\Delta_{n}e^{-i\phi_{n}}&\mu_{n}-V_{\mathrm{Z}}&0\\ \Delta_{n}e^{-i\phi_{n}}&0&0&\mu_{n}+V_{\mathrm{Z}}\end{pmatrix}_{\!n,n}$
$+$ $\sum_{n=1}^{N-1}t_{n}\begin{pmatrix}e^{i\boldsymbol{\lambda}_{n}\cdot\boldsymbol{\sigma}}&0\\ 0&-e^{-i\boldsymbol{\lambda}_{n}\cdot\boldsymbol{\sigma}^{T}}\end{pmatrix}_{\!n,n+1}+\mathrm{h.c.},$ (9)
with the onsite (dot) block $(\quad)_{n,n}$ and the offsite hopping (inter-dot) block $(\quad)_{n,n+1}$—both represented by $4\times 4$ matrices. For a system consisting of three quantum dots ($N=3$), the full BdG Hamiltonian $H^{\mathrm{BdG}}$ is a $12\times 12$ matrix.
Our goal is to simulate the Kitaev chain (KC) model, i.e., to tune the $H^{\mathrm{BdG}}$ (parameters) to mimic KC Hamiltonian form. The first step is that each QD should have a pair of zero energy levels $E_{n}=0$. Diagonalizing a (separated) single dot block $(\quad)_{n,n}$ yields the condition for this to happen: $V_{\mathrm{Z}}=\sqrt{\Delta^{2}_{n}+\mu^{2}_{n}}$ or equivalently $\Delta_{n}=\sqrt{V^{2}_{\mathrm{Z}}-\mu^{2}_{n}}$. When this condition is satisfied, each dot has a pair of fermionic excitations $a_{n},a^{\dagger}_{n}$, where
$a_{n}=\frac{1}{\sqrt{2V_{\mathrm{Z}}}}\left(e^{-\frac{i\phi_{n}}{2}}\sqrt{V_{\mathrm{Z}}+\mu_{n}}\,c^{\dagger}_{n\uparrow}-e^{\frac{i\phi_{n}}{2}}\sqrt{V_{\mathrm{Z}}-\mu_{n}}\,c_{n\downarrow}\right)\!.$ (10)
The energy of $a_{n}$ ($a^{\dagger}_{n}$) is zero. Two other modes are separated by the energy $\pm 2V_{\mathrm{Z}}$. Let’s now assume that the hopping is much smaller than the energy of the excited state, $t\ll 2V_{\mathrm{Z}}$, we may project the Hamiltonian (A) onto the Hilbert space spanned by a set $\{a_{1},a^{\dagger}_{1},a_{2},a^{\dagger}_{2},\dots,a_{N},a^{\dagger}_{N}\}$. To simplify the system definition, we take $\boldsymbol{\lambda}_{n}\cdot\boldsymbol{\sigma}=\lambda_{n}\sigma_{y}$, which follows from choosing the Rashba vector direction $\rho=\xi=\frac{\pi}{2}$. The resulting projected Hamiltonian is:
$\tilde{H}^{\mathrm{BdG}}=\sum_{n=1}^{N-1}\frac{t_{n}}{V_{\mathrm{Z}}}\begin{pmatrix}T_{n}&-D_{n}\\ D_{n}&-T^{\ast}_{n}\end{pmatrix}_{\!n,n+1}+\mathrm{h.c.},$ (11)
where $T_{n}=-(\mu_{n}\cos{\frac{\phi_{n+1}-\phi_{n}}{2}}-iV_{\mathrm{Z}}\sin{\frac{\phi_{n+1}-\phi_{n}}{2}})\cos{\lambda_{n}}$, and $D_{n}=-\Delta_{n}\cos{\frac{\phi_{n+1}-\phi_{n}}{2}}\sin{\lambda_{n}}$. The goal is to simulate the KC model, which, e.g. for $N=3$ takes the form (in the same Nambu representation $(a_{1},a^{\dagger}_{1},a_{2},a^{\dagger}_{2},a_{3},a^{\dagger}_{3})$ as $\tilde{H}^{\mathrm{BdG}}$):
$H^{\mathrm{KC}}=\begin{pmatrix}\mu_{1}&0&t^{\prime}_{1}&-\Delta^{\prime}_{1}&0&0\\ 0&-\mu_{1}&\Delta^{\prime}_{1}&-t^{\prime}_{1}&0&0\\ t^{\prime}_{1}&\Delta^{\prime}_{1}&\mu_{2}&0&t^{\prime}_{2}&-\Delta^{\prime}_{2}e^{i\Phi}\\ -\Delta^{\prime}_{1}&-t^{\prime}_{1}&0&-\mu_{2}&\Delta^{\prime}_{2}e^{i\Phi}&-t^{\prime}_{2}\\ 0&0&t^{\prime}_{2}&\Delta^{\prime}_{2}e^{-i\Phi}&\mu_{3}&0\\ 0&0&-\Delta^{\prime}_{2}e^{-i\Phi}&-t^{\prime}_{2}&0&-\mu_{3}\end{pmatrix}\!$. (12)
KC Hamitonian, Eq. 12, at the sweet-spot $\mu_{n}=0$, $t^{\prime}_{n}=\Delta^{\prime}_{n}$ possesses two MZMs. For the Rashba–Zeeman $s$-wave Hamiltonian $\tilde{H}^{\mathrm{BdG}}$, Eq. 11, to simulate $H^{\mathrm{KC}}$, Eq. 12, at the sweet-spot, the following conditions must be satisfied (assuming $\Phi=0$):
$\phi_{n+1}$ $=\phi_{n},$
$\mu_{n}\cos{\lambda_{n}}$ $=\Delta_{n}\sin{\lambda_{n}}.$ (13)
The latter yields $\lambda_{n}=\arctan(\frac{\mu_{n}}{\Delta_{n}})$ which, together with $\mu_{n}=\sqrt{V^{2}_{\mathrm{Z}}-\Delta^{2}_{n}}$, gives the approximate position of the parameters sweet-spot for the QDs Hamiltonian $H$. Both conditions are also discussed in the adaptive tuning proposal of Fulga et al. [14], as well as in the recent experimental realization of short KCs of Bordin et al. [4].
Appendix B Appendix B: Analyzing tuning procedure for hopping $t$ parameter
Figure 7: Three-QD chain Hamiltonian $H$ as a function of the offset $\mu$ added to local potentials: $\mu_{n}\rightarrow\mu_{n}+\mu-0.6$ meV (vertical dashed line marks the reference $\mu=0.6\,\mathrm{meV}$). Additionally (twin $y$-axis) $\mathcal{M}$ metric is plotted, and eigenvalues are colored (left column) with edge occupations $O_{edge}^{(i)}$, or (right column) electron-hole symmetry $S_{\mathrm{eh}}^{(i)}$ defined in the Methods section. Plots are presented for (a) reference parameters modified by setting $t_{0}=0.06\,\mathrm{meV}$ and $t_{1}=0.06\,\mathrm{meV}$; (b) parameters from (a) with PINNATuned $\{\mu_{n},t_{n},\lambda_{n}\}$; (c) parameters from (a) with PINNATuned $\{\mu_{n},V_{\mathrm{Z}}\}$.
Here we present an additional example of an autotuning procedure for the parameter $t$. In Fig. 7(a), we plot the $H$ eigenspectrum for detuned parameters $t_{0}=t_{1}=0.06$ meV, fixed to reduce $\mathcal{M}$ value at reference $\mu=0.6$ meV. It is visible that such a setting of $t_{0}$ and $t_{1}$ does not make zero modes vanish completely, but instead moves them to a lower $\mu$ regime. Simultaneously, the zero-energy states are far less gapped from the rest of the spectrum than for the reference $t=0.25$ meV setup.
In the next step we try to tune the parameters $\{\mu_{n},t_{n},\lambda_{n}\}$ using the NN model, as presented in Fig 7(b). We observe that PINNAT adjusts $\mu_{n}$ reducing the average $\langle\mu_{n}\rangle=0.51$, thereby shifting the energy offset and restoring the previously observed zero modes. Additionally, further setting $t_{0}=0.11$ meV, $t_{1}=0.1$ meV, $\lambda_{0}=0.38\,\pi$ and $\lambda_{1}=0.26\,\pi$ parameters allows NN to increase the $\mathcal{M}$ value and as a result, increase the topological gap – separation between zero modes and the bulk spectrum.
Finally, we can adjust the detuning in $t_{0}$ and $t_{1}$ using the second version of PINNAT model that sets $\{\mu_{n},V_{\mathrm{Z}}\}$ parameters – Fig. 7(c). Similarly like before, $\mu_{n}$ are shifted so that their average value is decreased: $\langle\mu_{n}\rangle=0.54$ meV, and zero modes emergence is observed. Moreover, the NN model proposes to set $V_{\mathrm{Z}}=0.24$ meV, yielding a higher value of $\mathcal{M}$ and an even larger topological gap than that in Fig. 7(b), where the model tunes $t_{n}$ and $\lambda_{n}$.
Nevertheless, similarly to the example presented in the main text, we can observe that both models are capable of correcting given parameters to produce $H$ with emergent zero-energy states, which are likely to maintain MZMs properties.
B.1 Appendix C: Scaling the system size
Figure 8: Iterative autotuning procedure for model adjusting $\{\mu_{n},V_{\mathrm{Z}}\}$ for system composed of 7 QDs. In the left plot, we present $\mathcal{M}$ map before tuning. Then, in the center, parameters are tuned with a single step of $N$ corrections. Finally, in the right plot, there is a map plotted for parameters tuned within 10 subsequent steps of $N$ corrections.
While analyzing results for a quantum chain composed of 3 QDs only proves the conceptual validity of the proposed approach, the natural question arises whether it is possible to repeat a similar procedure for larger systems. For this reason, we manually tune the Rashba-Zeeman-BCS Hamiltonian, Eq. 1, to realize MZMs for $N=7$ quantum dots. Such reference parameters include $\mu=0.35\,\mathrm{meV}$, $t=0.25\,\mathrm{meV}$, $\lambda=0.27\,\pi$, $\rho=\xi=\frac{\pi}{2}$, $V_{\mathrm{Z}}=0.5\,\mathrm{meV}$, and $\Delta=0.35\,\mathrm{meV}$. In contrast to 3-QDs systems, to train the PINNAT model, for the 7-QDs chain, we could not use entirely random parameter sampling due to a much larger parameter space and far less number of samples with $\mathcal{M}>0$. Instead, we generated 10 000 samples, randomly shifting the parameters from the reference values with the restriction that $\mathcal{M}$ value is larger than $\tau=0.7$. Subsequently, in each training epoch, the samples are independently randomized to decrease the value of $\mathcal{M}$. For half of the samples, $\mathcal{M}$ is reduced to 0, and for the remaining 5 000 samples, $\mathcal{M}$ is decreased to the interval $(0,\tau]$. Next, a set of 16 conductance maps is determined: to reduce the conductance tensor size we use only the $G_{\mathrm{LL}}$ and $G_{\mathrm{RR}}$ pair as a function of 8 parameters: $\mu_{n}$, for $n$ iterating over each dot, and $V_{\mathrm{Z}}$ (16 maps in total). Finally, the PINNAT model is trained to increase the value of $\mathcal{M}$ by adjusting the $\{\mu_{n},V_{\mathrm{Z}}\}$ parameters.
The results of training the model are shown in Fig. 8, where the $\mathcal{M}$ metric is plotted for the previously introduced 10-step autotuning procedure. In contrast to the 3-QDs system, Fig. 4(b), the PINNAT model trained to produce corrections for 7-QDs chain tends to tune the system to a lesser extent. However, the increase in $\mathcal{M}$ value is still visible even after a single step of corrections. Moreover, within the 10 steps, the PINNAT model demonstrates its ability to correct parameter detuning in the whole training distribution and also shows the potential of generalization beyond the training regime.
While these results prove that it is possible to scale the PINNAT model with an autotuning approach to larger systems, several limitations to the scaling should be highlighted. In particular, the dataset construction requires filtering Hamiltonians that exhibit emerging Majorana modes, while the training procedure involves repeated direct diagonalization within the loss evaluation (via the $\mathcal{M}$ metric) and the computation of conductance maps. As a result, both the data generation time and the training time scale approximately exponentially with the number of quantum dots in the system. This behavior is illustrated in Table 1, which summarizes the observed computational costs for systems of different sizes. While small systems can be handled almost instantaneously, the computational cost increases rapidly. For example, generating the dataset for a 10-dot system requires $\sim 17$ hours, and a single training epoch takes $\sim 5.5$ hours. Therefore, extending the method to larger systems will likely require additional optimizations, such as limiting the number of input conductance maps, reducing the number of tuned parameters, narrowing the parameter search ranges, or constructing a more time-effective target function.
Number of dots Data generation time [h] Training time per epoch [h]
3 online 0.28
5 2 0.42
7 5 1.17
10 17 5.50
Table 1: Observed scaling of data generation and model training times with system size.
Additionally, although the eigendecomposition used in the loss evaluation is differentiable almost everywhere, the gradients associated with eigenvectors can become numerically ill-conditioned near degeneracies or level crossings. This effect may introduce noise in the backpropagated gradients and can potentially limit the scalability of eigenvalue-based learning procedures. Various approaches have been proposed to mitigate this issue in differentiable eigensolvers [54, 25]. However, in our experiments, we observed stable and well-behaved training for systems up to 10 quantum dots (corresponding to $40\times 40$ Hamiltonian matrices), with no noticeable degradation of the optimization process.
B.2 Appendix D: Comparison of tuned parameters with Majorana sweet-spot conditions
Table 2: Comparison of reference and PINNAT parameter sets for the three-dot chain. Energies are given in meV. In the presented sets $\Delta_{n}$ does not change with $n$: $\Delta_{n}=\Delta$.
Configuration $\langle\mu_{n}\rangle$ $V_{\mathrm{Z}}$ $\Delta$ $t_{1}$ $t_{2}$ $\lambda_{1}$ $\lambda_{2}$ $\frac{V_{\mathrm{Z}}}{\max_{n}{t_{n}}}$ $\frac{V_{\mathrm{Z}}}{\sqrt{\Delta^{2}+\langle\mu_{n}\rangle^{2}}}$ $\frac{\langle\lambda_{n}\rangle}{\arctan\left(\frac{\langle\mu_{n}\rangle}{\Delta}\right)}$
Reference parameters 0.60 0.50 0.25 0.25 0.25 $0.27\pi$ $0.27\pi$ 2.0 0.77 0.72
PINNAT $\{\mu_{n},t_{n},\lambda_{n}\}$ (Fig. 5c) 0.54 0.50 0.25 0.20 0.25 $0.67\pi$ $0.41\pi$ 2.0 0.83 1.49
PINNAT $\{\mu_{n},V_{\mathrm{Z}}\}$ (Fig. 5d) 0.58 0.31 0.25 0.25 0.25 $0.62\pi$ $0.41\pi$ 1.24 0.49 1.39
PINNAT $\{\mu_{n},t_{n},\lambda_{n}\}$ (Fig. 7b) 0.51 0.50 0.25 0.11 0.10 $0.38\pi$ $0.26\pi$ 4.55 0.88 0.90
PINNAT $\{\mu_{n},V_{\mathrm{Z}}\}$ (Fig. 7c) 0.54 0.24 0.25 0.06 0.06 $0.27\pi$ $0.27\pi$ 4.0 0.4 0.75
To facilitate comparison with previously proposed parameter regimes for Majorana zero modes in few-site Kitaev chains, Table 2 lists representative sets of PINNAT parameters together with the reference values used in the simulations. In addition to the raw parameters, dimensionless ratios are reported that quantify proximity to the sweet-spot relations discussed in Appendix A.
The PINNAT configurations consistently fall within the expected Majorana-supporting regime. In particular, the Zeeman energy exceeds the inter-dot couplings in all cases, as indicated by $V_{\mathrm{Z}}/\max_{n}t_{n}>1$, often by a significant margin. This places the system in a regime where Zeeman splitting dominates over inter-dot hybridization, which is favorable for the emergence of topological phases.
Furthermore, the ratio $V_{\mathrm{Z}}/\sqrt{\Delta^{2}+\langle\mu_{n}\rangle^{2}}$ remains of order unity, indicating that the condition $V_{\mathrm{Z}}=\sqrt{\Delta^{2}+\mu_{n}^{2}}$ for having a single fermionic pair at zero energy per dot is approximately satisfied. Finally, the relation between the Rashba spin-orbit length and the onsite potential, captured by $\langle\lambda_{n}\rangle/\arctan\frac{\langle\mu_{n}\rangle}{\Delta}$, also remains of order unity across the considered configurations, indicating consistency with the analytical condition derived in Appendix A.
Overall, while the analytical sweet-spot conditions are not exactly fulfilled, the PINNAT parameters systematically lie in their approximate regime of validity, demonstrating that the autotuning procedure identifies physically consistent Majorana-supporting configurations.